# An overview of condensation phenomenon in deep learning

### v1

@@ 1 Introduction
One such phenomenon is the over-parameterization puzzle, which has led to a deeper understanding of neural network generalization [Bre95, ZBH+17]. This puzzle reveals that a neural network can generalize well even when the number of parameters far exceeds the number of training data points. This observation challenges traditional learning theory, which typically improves generalization by imposing constraints on model complexity [Vap13]. In contrast, the generalization of large neural networks appears to be largely independent of superficial complexity, such as the number of parameters. Instead, the optimization trajectory plays a crucial role in locating a minimum with specific properties among various types of minima. Empirical studies have shown that smaller batch sizes in Stochastic Gradient Descent (SGD) tend to lead to flatter minima, which is associated with better generalization [KMN+16]. This led to the development of sharpness-aware minimization (SAM) [FKMN21] techniques that further improve generalization performance. Additionally, recent works have shown that the noise covariance induced by SGD aligns with the Hessian of the loss landscape [ZWY+18, WME18, FT21], providing further insights into the optimization dynamics.
@@
To further investigate the simplicity bias, several studies have analyzed the evolution of neural network parameters during training. Two distinct regimes [LXMZ21, ZZJ+22] have been identified: the linear regime, in which parameters initialized with relatively large values undergo minimal changes during training, and the nonlinear regime, where smaller initializations result in more substantial parameter adjustments [RVE18, CB18]. In the linear regime, the behavior of the neural network closely resembles that of kernel methods, with the neural tangent kernel (NTK) [JGH18, COB19] being a prominent example. The transition between the linear and nonlinear regimes represents a critical phase, with mean-field dynamics being a typical example [MMM19, SS20, RVE18]. It is in the nonlinear regime that a universal condensation phenomenon occurs [LXMZ21, ZZL+22, ZZJ+22]. In this paper, we aim to overview previous works on this phenomenon and provide a unified description of condensation:
@@
This condensed regime represents a state in which neurons in the same layer condense into a few distinct groups, with neurons within each group performing similar functions. This clustering phenomenon implies that a wide neural network can behave similarly to a much narrower network. Early in the nonlinear training process, neurons in the same layer tend to group into a small number of clusters [MBG18, PL21, LLWA21, BPVF22, ZZL+22, ZZJ+22, MMV24, WM24]. As training progresses, the number of clusters increases, which facilitates fitting. Thus, the condensation phenomenon offers a mechanism for the increasing complexity of the network’s outputs as training progresses.
In this paper, we present experiments with various neural network architectures to demonstrate the ubiquity of the condensation phenomenon in nonlinear training [LMW+21, ZZL+22, ZZLX23]. We also explore how dropout [SHK+14] implicitly induces a bias toward condensation [ZX24, ZLLX24]. Furthermore, we examine the origins of condensation from the perspectives of loss landscapes and training dynamics. The condensation phenomenon suggests a potential pruning strategy, where network size can be reduced without sacrificing generalization ability [ZZLX21, CX24]. This insight also leads to a novel optimistic estimation of the sample size required to recover a target function based on a perfectly condensed network [ZZZ+23] rather than relying on superficial network complexity, where the latter often leads to overly conservative estimates.
Moreover, the condensation phenomenon, originally observed in simple two-layer neural networks, provides a deeper understanding of the reasoning and memorization processes in transformer models, particularly for language tasks [ZLW+24, ZLW+25]. This understanding could also inform methods for training transformer networks with improved reasoning capabilities.
@@
This phenomenon has been characterized in various ways throughout the literature. [MBG18] described it as a quantization effect where weight vectors tend to concentrate in finite directions due to gradient descent. [BG19] referred to it as the weight clustering effect. [COB19] provided an illustrative example of non-lazy training. [PL21] named this behavior a form of inductive bias. Several works investigated this behavior of neurons within the same layer and named it “alignment/get align”[JT19, LLWA21, BPVF22, CEL23, MMV24, BF24]. [KH24a, KH24b] termed this phenomenon “directional convergence”.
@@ 2 Concept of condensation
Two vectors have the same (or opposite) directions when their cosine similarity $D(\bm{u}_{1},\bm{u}_{2})$ is $1$ (or $-1$).
@@ 3 Condensation process during the training
where $\sigma(x)=\mathrm{ReLU}(x)$. The width of the hidden layer is $m=100$, and the learning rate is $0.1$. The parameters are initialized by $\mathcal{N}(0,\frac{1}{m^{4}})$. The training data is evenly sampled in $[-1,1]$.
The features $\{(\theta_{k},A_{k})\}_{k}$ during the training process are shown in Fig. 1. We observe that, as training progresses, the neurons in the network condense into a few isolated orientations, and the number of these orientations increases. A similar training process is shown in [BF24] on a piece-wise linear target function proposed in [SBBV23].
@@
(a) epoch $=100$
(b) epoch $=1000$
(c) epoch $=5000$
(d) epoch $=10000$
(e) epoch $=12000$
(f) epoch $=100000$
@@ 4.1 Condensation in the synthetic data
Consider a target function $f(x)=\mathrm{Tanh}(x)$. We also use a two-layer Tanh NN to fit the target function. The width of the hidden layer is $m=1000$, and the learning rate is $0.03$. The training data is evenly sampled in $[-15,15]$. The parameters are also initialized by $\mathcal{N}(0,\frac{1}{m^{\gamma}})$.
@@ 4.2 Condensation in the CNNs
Figure 3: Small initialization (convolutional and fully connected layers initially follow $\mathcal{N}(0,96^{-8})$) for single-layer CNN training in its final stage of convergence. The activation function is $\rm{tanh}(x)$. If neurons are in the same dark blue block, then $D(\bm{u},\bm{v})\sim 1$ (in beige blocks, $D(\bm{u},\bm{v})\sim-1$), indicating that their input weight directions are the same (opposite). Colors represent $D(\bm{u},\bm{v})$ of two convolution kernels, with indices shown on the horizontal and vertical axes respectively. The training set is MNIST. The output layer uses softmax, the loss function is cross-entropy, and the optimizer is Adam with full batch training. Convolution kernel size $m=3$, learning rate $=2\times 10^{-4}$. Training continues until $100\%$ accuracy is achieved on the training set, at this point, the test set accuracy is $97.62\%$.
@@ 5 Phase diagram: when condensation happens
Empirically, we have found that in non-linear training regime, condensation is a very common phenomenon. In Ref. [LXMZ21], to characterize the non-linear and linear regimes, we consider a two-layer NN with $m$ hidden neurons
@@
Theorem 1 (Informal statement [LXMZ21]).
@@
Theorem 2 (Informal statement [LXMZ21]).
@@
For the non-linear regime, we find that condensation is a unique feature, therefore, we name it condensation regime. For three-layer ReLU neural networks, we found similar phase diagrams for the dynamics of each layer [ZZJ+22].
The study of phase diagrams provides valuable insights into how to appropriately tune parameter initialization when scaling up network sizes. A commonly used initialization method involves sampling the parameters from a Gaussian distribution $\mathcal{N}(0,\frac{1}{m^{\gamma}})$ (or a uniform $[-\frac{1}{m^{\gamma}},\frac{1}{m^{\gamma}}]$ distribution), where $m$ represents the input dimension or the average of the input and output dimensions. When scaling up network sizes, to maintain similar dynamic behavior, it is crucial not to fix the initialization variance, but rather to keep $\gamma$ fixed.
Figure 5: Phase diagram of two-layer ReLU NNs at infinite-width limit. The marked examples are studied in existing literature. Table is from Ref. [LXMZ21].
@@ 6.1 Initial condensation
Experiments in [ZZL+22] suggest that the maximum number of condensed directions for input weights is no greater than $2p$. Additionally, theoretical analysis is provided for the case of $p=1$, as well as for any $p$ with one-dimensional input. For the case of $p=1$, [CLL+23] further estimates the time required for initial condensation in two-layer NNs. The following example illustrates how the activation function can influence the initial condensed directions. As is shown in Fig. 6, when employing $\mathrm{Tanh}$ as the activation, there are a pair of opposite condensed directions. When the activation function is $\mathrm{xTanh}$, there are two pairs of opposite condensed directions.
In the case of $p=1$, several works investigate different scenarios. [CL24] shows that three layer NNs will have condensed solutions at the initial stage with some assumptions. [ZZLX23] analyzes the initial condensation of two-layer convolutional NNs. [CLW24] analyzes the subsequent loss descent and the second loss plateau after the initial condensation stage.
@@
For the second case, [MBG18] shows that in the limit of infinitesimal initial weights and learning rate, two-layer ReLU NN will first align at a discrete set of possible directions before the loss descent. [PL21] analyzes a more concrete setting on the orthogonally separable data and the neurons will asymptotically converge to two neurons: the positive max-margin vector and the negative max-margin vector. [BPVF22] investigates the time of the early alignment stage when the data forms an orthonormal family. [CEL23] observes that when using a two layer ReLU network to learn a target function of one neuron with correlated inputs, the neurons will first align and will not separate during training. [WM24] estimates the time of the early alignment phase in the binary classification problem of effectively two data points, which are separated by small angles, and [MMV24] looses the data assumption to that the data are positively correlated when they have the same labels. [BF24] demonstrates a quantitative analysis of the initial condensation of both regression and classification and general datasets in two layer NNs. They also give an example that the initial condensation will do harm to the final convergence with the initialization that $|a_{j}|\geq||w_{j}||$. [KH24a, KH24b] extends the analysis of early alignment to homogeneous neural networks, with [KH24a] exploring alignment dynamics that near saddle points beyond initialization on two-homogeneous NNs. [LLWA21] demonstrates that a two-layer leaky ReLU NN with linear separable and symmetric data will align in the first phase and finally reach a global-max-margin linear classifier.
@@ 6.2 Embedding principle
To understand the similarities and differences among networks with varying widths, [ZZLX21] introduced an embedding principle, which states that the loss landscape of any neural network “contains” all critical points of all narrower networks. Similar ideas are also studied in [FA00, FYMT19, SGJ+21]. Specifically, for a narrow network at a critical point, if a neuron is split into two neurons in the following manner: the new neurons have the same input weights as the original one, and the sum of the output weights of the two new neurons to a subsequent neuron equals the output weight of the original neuron to that subsequent neuron, then the wider network will also be at a critical point. This explains the similarities shared by networks of various widths. It is important to note that the wider network can be regarded as in a condensed state.
@@ 7.2 Reasoning ability of Transformer
[ZLW+24, ZLW+25] explore the role of condensation in enhancing the reasoning ability of Transformer models. The task is to study a composite function composed of several simple functions, i.e., addition and subtraction. Specifically, we define 4 simple functions (denoted as function 1, 2, 3, 4) and they can form 16 composite functions. We use 14 composite functions for training and leave the composition of functions 3 and 4 for testing (i.e., (3, 4) and (4, 3)). In distribution (ID) generalization refers to the accuracy of training composite functions with unseen computed numbers, while out of distribution (OOD) refers to the accuracy of test composite functions.
The parameters of the transformer network are initialized by $\mathcal{N}(0,\frac{1}{m^{\gamma}})$, where $m$ is the width of the layer. We observe that as the initialization rate $\gamma$ increases, i.e., initialization scale decreases, the transformer network learns the data respectively by the following four patterns: i) The network only remembers training data, and shows no generalization on any test data of seen or unseen composite functions; ii) The network can generalize to the seen composite function operating on unseen numbers, but not on the solution of unseen composite function (3, 4) or (4, 3), in addition, the network output of composite function (3, 4) and (4, 3) shows no symmetry; iii) Similar to (ii) but the network output of composite function (3, 4) and (4, 3) is symmetric; iv) The network generalizes to all composite functions, which indicates the network learns all primitive functions. This simple experiment shows that $\gamma$ can well tune the network to bias towards memorizing or reasoning data. Additionally, as shown in Fig. 10, we notice that during this process, the phenomenon of condensation becomes increasingly pronounced, suggesting a strong correlation between the condensation phenomenon and the model’s reasoning ability. A straightforward rationale is as follows: since the network strongly favors condensation, it tends to learn the data with the lowest possible complexity. Clearly, if the model can uncover the underlying simple function, it only needs to memorize a few simple functions rather than numerous data pairs. Consequently, it can explain the data with minimal effective complexity. An analysis of the initial training stage for reasoning bias of language models with small initialization further enhances the relation between condensation and reasoning [YZX25].

### v2

@@ 1 Introduction
One such phenomenon is the over-parameterization puzzle, which has led to a deeper understanding of neural network generalization [BRE95, ZBH+17]. This puzzle reveals that a neural network can generalize well even when the number of parameters far exceeds the number of training data points. This observation challenges traditional learning theory, which typically improves generalization by imposing constraints on model complexity [VAP13]. In contrast, the generalization of large neural networks appears to be largely independent of superficial complexity, such as the number of parameters. Instead, the optimization trajectory plays a crucial role in locating a minimum with specific properties among various types of minima. Empirical studies have shown that smaller batch sizes in Stochastic Gradient Descent (SGD) tend to lead to flatter minima, which is associated with better generalization [KMN+16]. This led to the development of sharpness-aware minimization (SAM) [FKM+21] techniques that further improve generalization performance. Additionally, recent works have shown that the noise covariance induced by SGD aligns with the Hessian of the loss landscape [ZWY+18, WME18, FT21], providing further insights into the optimization dynamics.
@@
To further investigate the simplicity bias, several studies have analyzed the evolution of neural network parameters during training. Two distinct regimes [LXM+21b, ZZJ+22a] have been identified: the linear regime, in which parameters initialized with relatively large values undergo minimal changes during training, and the nonlinear regime, where smaller initializations result in more substantial parameter adjustments [RV18, CB18]. In the linear regime, the behavior of the neural network closely resembles that of kernel methods, with the neural tangent kernel (NTK) [JGH18, COB19] being a prominent example. The transition between the linear and nonlinear regimes represents a critical phase, with mean-field dynamics being a typical example [MMM19, SS20, RV18]. It is in the nonlinear regime that a universal condensation phenomenon occurs [LXM+21b, ZZL+22b, ZZJ+22a]. In this paper, we aim to overview previous works on this phenomenon and provide a unified description of condensation:
@@
This condensed regime represents a state in which neurons in the same layer condense into a few distinct groups, with neurons within each group performing similar functions. This clustering phenomenon implies that a wide neural network can behave similarly to a much narrower network. Early in the nonlinear training process, neurons in the same layer tend to group into a small number of clusters [MBG18, PL21, LLW+21, BPF22, ZZL+22b, ZZJ+22a, MMV24, WM24]. As training progresses, the number of clusters increases, which facilitates fitting. Thus, the condensation phenomenon offers a mechanism for the increasing complexity of the network’s outputs as training progresses.
In this paper, we present experiments with various neural network architectures to demonstrate the ubiquity of the condensation phenomenon in nonlinear training [LMW+21a, ZZL+22b, ZZL+23]. We also explore how dropout [SHK+14] implicitly induces a bias toward condensation [ZX24, ZLL+24a]. Furthermore, we examine the origins of condensation from the perspectives of loss landscapes and training dynamics. The condensation phenomenon suggests a potential pruning strategy, where network size can be reduced without sacrificing generalization ability [ZZL+21, CX24]. This insight also leads to a novel optimistic estimation of the sample size required to recover a target function based on a perfectly condensed network [ZZZ+23] rather than relying on superficial network complexity, where the latter often leads to overly conservative estimates.
Moreover, the condensation phenomenon, originally observed in simple two-layer neural networks, provides a deeper understanding of the reasoning and memorization processes in transformer models, particularly for language tasks [ZLW+24b, ZLW+25]. This understanding could also inform methods for training transformer networks with improved reasoning capabilities.
@@
This phenomenon has been characterized in various ways throughout the literature. [MBG18] described it as a quantization effect where weight vectors tend to concentrate in finite directions due to gradient descent. [BG19] referred to it as the weight clustering effect. [COB19] provided an illustrative example of non-lazy training. [PL21] named this behavior a form of inductive bias. Several works investigated this behavior of neurons within the same layer and named it “alignment/get align”[JT19, LLW+21, BPF22, CEL23, MMV24, BF24]. [KH24a, KH24b] termed this phenomenon “directional convergence”.
@@ 2 Concept of condensation
Two vectors have the same (or opposite) directions when their cosine similarity $D(\bm{u}_{1},\bm{u}_{2})$ is $1$ (or 1-1).
@@ 3 Condensation process during the training
where $\sigma(x)=\mathrm{ReLU}(x)$. The width of the hidden layer is $m=100$, and the learning rate is $0.1$. The parameters are initialized by $\mathcal{N}(0,\frac{1}{m^{4}})$. The training data is evenly sampled in [1,1][-1,1].
The features $\{(\theta_{k},A_{k})\}_{k}$ during the training process are shown in Fig. 1. We observe that, as training progresses, the neurons in the network condense into a few isolated orientations, and the number of these orientations increases. A similar training process is shown in [BF24] on a piece-wise linear target function proposed in [SBB+23].
@@
[YZZ24] use an effective rank to measure the linear dependence among different neurons and found a staircase phenomenon, that is, the effective rank gradually increases in the training process, similar to a staircase curve. The staircase phenomenon is closely related to the condensation process. As the neurons condense into more and more groups, neurons in different groups are linearly independent, leading to the increase of effective rank. However, [YZZ24] do not notice that the scaling of initialization is a critical index to such non-linear phenomena, that is, the phase diagram of different training dynamics induced by initialization studied in [LXM+21b].
(a) epoch =100=100
(b) epoch =1000=1000
(c) epoch =5000=5000
(d) epoch =10000=10000
(e) epoch =12000=12000
(f) epoch =100000=100000
@@ 4.1 Condensation in the synthetic data
Consider a target function $f(x)=\mathrm{Tanh}(x)$. We also use a two-layer Tanh NN to fit the target function. The width of the hidden layer is $m=1000$, and the learning rate is $0.03$. The training data is evenly sampled in [15,15][-15,15]. The parameters are also initialized by $\mathcal{N}(0,(\frac{1}{m^{\gamma}})^{2})$, where $\frac{1}{m^{\gamma}}$ is the standard deviation.
@@ 4.2 Condensation in the CNNs
Figure 3: Small initialization (convolutional and fully connected layers initially follow $\mathcal{N}(0,96^{-8})$) for single-layer CNN training in its final stage of convergence. The activation function is $\rm{tanh}(x)$. If neurons are in the same dark blue block, then $D(\bm{u},\bm{v})\sim 1$ (in beige blocks, $D(\bm{u},\bm{v})\sim-1$), indicating that their input weight directions are the same (opposite). Colors represent $D(\bm{u},\bm{v})$ of two convolution kernels, with indices shown on the horizontal and vertical axes respectively. The training set is MNIST. The output layer uses softmax, the loss function is cross-entropy, and the optimizer is Adam with full batch training. Convolution kernel size $m=3$, learning rate $2\times 10^{-4}$. Training continues until $100\%$ accuracy is achieved on the training set, at this point, the test set accuracy is $97.62\%$.
@@ 5 Phase diagram: when condensation happens
Empirically, we have found that in non-linear training regime, condensation is a very common phenomenon. In Ref. [LXM+21b], to characterize the non-linear and linear regimes, we consider a two-layer NN with $m$ hidden neurons
@@
Theorem 1 (Informal statement [LXM+21b]).
@@
Theorem 2 (Informal statement [LXM+21b]).
@@
For the non-linear regime, we find that condensation is a unique feature, therefore, we name it condensation regime. For three-layer ReLU neural networks, we found similar phase diagrams for the dynamics of each layer [ZZJ+22a].
The study of phase diagrams provides valuable insights into how to appropriately tune parameter initialization when scaling up network sizes, which we elaborate in the next section.
Figure 5: Phase diagram of two-layer ReLU NNs at infinite-width limit. The marked examples are studied in existing literature. Table is from Ref. [LXM+21b].
6 Initialization scaling exponent as a hyperparameter
The study of phase diagrams provides valuable insights into how to appropriately tune parameter initialization when scaling up network sizes. A commonly used initialization method involves sampling the parameters from a Gaussian distribution $\mathcal{N}(0,(\frac{1}{m^{\gamma}})^{2})$ (or a uniform $\frac{1}{m^{\gamma}},\frac{1}{m^{\gamma}}]$ distribution), where $\frac{1}{m^{\gamma}}$ is the standard deviation and $m$ represents the input dimension or the average of the input and output dimensions. When scaling up network sizes, to maintain similar dynamic behavior, it is crucial not to fix the initialization standard deviation, but rather to keep $\gamma$ fixed. The exponent $\gamma$ directly determines the training regime — whether the network operates in the linear regime or the condensation regime — and thus fundamentally shapes the optimization landscape and generalization behavior.
In current mainstream deep learning frameworks such as PyTorch, TensorFlow, JAX, and Megatron-LM, the initialization of network parameters is typically specified through named schemes. For example, PyTorch provides kaiming_uniform, kaiming_normal, xavier_uniform, and xavier_normal, each corresponding to a specific choice of the scaling exponent $\gamma$ (e.g., $\gamma=1/2$ for Kaiming and Xavier family). While these built-in schemes are convenient, they implicitly fix $\gamma$ and offer no straightforward mechanism for users to explore alternative values. In practice, switching from, say, $\gamma=1/2$ to $\gamma=3/4$ requires manually computing the desired standard deviation and calling low-level initialization functions, which is error-prone and obscures the underlying intent.
We propose that the initialization scaling exponent $\gamma$ should be explicitly exposed as a first-class hyperparameter in deep learning frameworks. Just as learning rate, weight decay, and momentum are standard tunable parameters in an optimizer, $\gamma$ governs a qualitatively different and equally important aspect of training: the dynamical regime of the network. A practical interface could take the form of an additional argument in the initialization API, e.g., nn.init.kaiming_normal_(tensor, gamma=0.5), or a global configuration option that adjusts all layer initializations consistently.
There are several concrete benefits to this proposal. First, it enables reproducible exploration of training regimes. Researchers and practitioners can systematically sweep over $\gamma$ values to identify whether their architecture and data benefit from condensation or linear dynamics, without rewriting initialization code. Second, it ensures consistency when scaling model sizes. As models grow wider or deeper, the phase diagram tells us that maintaining the same $\gamma$ preserves the dynamical regime, whereas naively preserving the standard deviation (a fixed $\sigma$) can inadvertently shift the network into a different regime, leading to unexpected training failures. Third, framework-level support for $\gamma$ facilitates the transfer of theoretical insights into practice. The phase diagram framework provides rigorous guidance on how $\gamma$ affects training, but this guidance can only be widely adopted if the relevant parameter is easy to control in code.
We encourage the deep learning community and framework developers to consider incorporating $\gamma$ as an explicit, tunable hyperparameter. This small interface change would bridge the gap between the theoretical understanding of training dynamics and everyday deep learning practice, enabling more principled initialization strategies as models continue to scale.
@@ 7.1 Initial condensation
Experiments in [ZZL+22b] suggest that the maximum number of condensed directions for input weights is no greater than $2p$. Additionally, theoretical analysis is provided for the case of $p=1$, as well as for any $p$ with one-dimensional input. For the case of $p=1$, [CLL+23] further estimates the time required for initial condensation in two-layer NNs. The following example illustrates how the activation function can influence the initial condensed directions. As is shown in Fig. 6, when employing $\mathrm{Tanh}$ as the activation, there are a pair of opposite condensed directions. When the activation function is $\mathrm{xTanh}$, there are two pairs of opposite condensed directions.
In the case of $p=1$, several works investigate different scenarios. [CL24] shows that three layer NNs will have condensed solutions at the initial stage with some assumptions. [ZZL+23] analyzes the initial condensation of two-layer convolutional NNs. [CLW24] analyzes the subsequent loss descent and the second loss plateau after the initial condensation stage.
@@
[…]
@@ 7.2 Embedding principle
To understand the similarities and differences among networks with varying widths, [ZZL+21] introduced an embedding principle, which states that the loss landscape of any neural network “contains” all critical points of all narrower networks. Similar ideas are also studied in [FA00, FYM+19, SGJ+21]. Specifically, for a narrow network at a critical point, if a neuron is split into two neurons in the following manner: the new neurons have the same input weights as the original one, and the sum of the output weights of the two new neurons to a subsequent neuron equals the output weight of the original neuron to that subsequent neuron, then the wider network will also be at a critical point. This explains the similarities shared by networks of various widths. It is important to note that the wider network can be regarded as in a condensed state.
@@ 8.2 Reasoning ability of Transformer
[ZLW+24b, ZLW+25] explore the role of condensation in enhancing the reasoning ability of Transformer models. The task is to study a composite function composed of several simple functions, i.e., addition and subtraction. Specifically, we define 4 simple functions (denoted as function 1, 2, 3, 4) and they can form 16 composite functions. We use 14 composite functions for training and leave the composition of functions 3 and 4 for testing (i.e., (3, 4) and (4, 3)). In distribution (ID) generalization refers to the accuracy of training composite functions with unseen computed numbers, while out of distribution (OOD) refers to the accuracy of test composite functions.
The parameters of the transformer network are initialized by $\mathcal{N}(0,(\frac{1}{m^{\gamma}})^{2})$, where $\frac{1}{m^{\gamma}}$ is the standard deviation and $m$ is the width of the layer. We observe that as the initialization rate $\gamma$ increases, i.e., initialization scale decreases, the transformer network learns the data respectively by the following four patterns: i) The network only remembers training data, and shows no generalization on any test data of seen or unseen composite functions; ii) The network can generalize to the seen composite function operating on unseen numbers, but not on the solution of unseen composite function (3, 4) or (4, 3), in addition, the network output of composite function (3, 4) and (4, 3) shows no symmetry; iii) Similar to (ii) but the network output of composite function (3, 4) and (4, 3) is symmetric; iv) The network generalizes to all composite functions, which indicates the network learns all primitive functions. This simple experiment shows that $\gamma$ can well tune the network to bias towards memorizing or reasoning data. Additionally, as shown in Fig. 10, we notice that during this process, the phenomenon of condensation becomes increasingly pronounced, suggesting a strong correlation between the condensation phenomenon and the model’s reasoning ability. A straightforward rationale is as follows: since the network strongly favors condensation, it tends to learn the data with the lowest possible complexity. Clearly, if the model can uncover the underlying simple function, it only needs to memorize a few simple functions rather than numerous data pairs. Consequently, it can explain the data with minimal effective complexity. An analysis of the initial training stage for reasoning bias of language models with small initialization further enhances the relation between condensation and reasoning [YZX25].